OpenStudy (anonymous):

If an object is dropped from a height of 144 feet, the function h(t)=-16t^2+144 gives the height of the object after t seconds. When will the object hit the ground?

myininaya (myininaya):

what is the height of the object from the ground if the object is on the ground?

OpenStudy (anonymous):

0.

OpenStudy (mertsj):

If the object is on the ground? I think it is 0 feet above the ground.

myininaya (myininaya):

right! h=0

myininaya (myininaya):

\[0=-16t^2+144\]

myininaya (myininaya):

So t is time t is the when that you are solving for

OpenStudy (anonymous):

how do you solve for t?

myininaya (myininaya):

\[0=-a^2+b\] If we wanted to solve this for a \[\text{I would add }a^2 \text{ on both sides}\] \[a^2=b\] Then I would take square root of both sides \[a=\pm \sqrt{b}\]

myininaya (myininaya):

So for your equation what would you add to both sides?

OpenStudy (anonymous):

9 seconds?

myininaya (myininaya):

\[9=t^2 ?\]

myininaya (myininaya):

is that what you are saying?

myininaya (myininaya):

you didn't do anything with the square part yet have you?

OpenStudy (anonymous):

\[\approx\]0.9083, -9.9083

myininaya (myininaya):

\[h(t)=-16t^2+144\] This is the function right? Now we want to know what t is when h is 0 \[0=-16t^2+144\] \[\text{ add } 16t^2 \text{ on both sides } \] \[16t^2=144 \]

myininaya (myininaya):

Divide both sides by 16 to undo the multiplication of 16 (Or in other words multiply both sides by 1/16)

OpenStudy (anonymous):

so 9 is correct.

myininaya (myininaya):

nope

myininaya (myininaya):

\[t^2=\frac{144}{16}\] \[t^2=9\]

myininaya (myininaya):

Do you know what to do from here?